OpenStudy (anonymous):

Evaluate -4(a+b) - 10a/b when a=8 and b=5

OpenStudy (anonymous):

-68

OpenStudy (anonymous):

How do you get the answer

OpenStudy (anonymous):

substitute the a and b values into the equation you've been provided.

OpenStudy (anonymous):

just beat ittttttttttttttttttttttttttttttttttt!

OpenStudy (anonymous):

i kno i did but after u enter it

OpenStudy (anonymous):

simple multiplication and division from there.

OpenStudy (anonymous):

show how u got that answer tho

OpenStudy (anonymous):

-4 (13)-10 8/5 then

OpenStudy (anonymous):

-4 * 13 = 52 \[\frac{ 10 * 8 }{ 5 }\] = 16 52-16=36

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I take that back, sorry the first line is incorrect. it should be -4*8 + -4*5 which equals -52

OpenStudy (anonymous):

then -52-16 =-68

OpenStudy (anonymous):

k